OpenStudy (anonymous):

The measures of each exterior angle in a regular hexagon is (4x + 15)°. Find the value of x. Show equations and all work that leads to your answer.

OpenStudy (anonymous):

measure of each exterior angle in a regular polygon= 360/n n= no of sides use that

OpenStudy (anonymous):

(4x + 15)°=360/6